An experiment was planned to compare the mean time (in days) required to recover from a common cold for persons given a daily dose of 4 mg of vitamin C, μ2, versus those who were not given a vitamin supplement, μ1. Suppose that 34 adults were randomly selected for each treatment category and that the mean recovery times and standard deviations for the two groups were as follows.

     No Vitamin
Supplement
4 mg
Vitamin C
Sample size 34 34
Sample mean     6.1     5.1
Sample standard deviation     2.4     1.9

(a) Suppose your research objective is to show that the use of vitamin C reduces the mean time required to recover from a common cold and its complications. Give the null and alternative hypotheses for the test.

H0: (μ1μ2) = 0 versus Ha: (μ1μ2) ≠ 0

H0: (μ1μ2) < 0 versus Ha: (μ1μ2) > 0   

H0: (μ1μ2) = 0 versus Ha: (μ1μ2) > 0

H0: (μ1μ2) ≠ 0 versus Ha: (μ1μ2) = 0

H0: (μ1μ2) = 0 versus Ha: (μ1μ2) < 0

Is this a one- or a two-tailed test?


(b) Conduct the statistical test of the null hypothesis in part (a) and state your conclusion. Test using

α = 0.05.

(Round your answers to two decimal places.)
z =

H0 is rejected. There is insufficient evidence to indicate that Vitamin C reduces the mean recovery time.H0 is not rejected. There is sufficient evidence to indicate that Vitamin C reduces the mean recovery time.    H0 is not rejected. There is insufficient evidence to indicate that Vitamin C reduces the mean recovery time.H0 is rejected. There is sufficient evidence to indicate that Vitamin C reduces the mean recovery time.
Answer #1

For No vitamin supplement :

x̅1 = 6.1, σ1 = 2.4, n1 = 34

For 4 mg Vitamin C :

x̅2 = 5.1, σ2 = 1.9, n2 = 34

a) Null and alternative hypotheses for the test.

H0: (μ1 − μ2) = 0 versus Ha: (μ1 − μ2) > 0

b) Test statistic:

z = (x̅1 - x̅2)/√(σ1²/n1 + σ2²/n2 ) = (6.1 - 5.1)/√(2.4²/34 + 1.9²/34) = 1.90

p-value :

p-value =1- NORM.S.DIST(1.9049, 1) = 0.0284

H0 is rejected. There is sufficient evidence to indicate that Vitamin C reduces the mean recovery time.